OpenStudy (saifoo.khan):

Binomial:

OpenStudy (saifoo.khan):

\[(2-5x)^6 (1+x/)^10 \]

OpenStudy (saifoo.khan):

its ^10.

OpenStudy (anonymous):

what do you want to do, simplify?

OpenStudy (saifoo.khan):

Find the coefficient of x in the expansion

OpenStudy (saifoo.khan):

plz chk q6 b.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i have no clue

myininaya (myininaya):

some how we need to expand with out really expanding it because thats to long to do like we can find the coefficient of x in 5x^2-2x+4 the coefficient of x is -2 so we got to see if there is some kind of trick we can do to finding the coefficient in your expansion

OpenStudy (saifoo.khan):

no, but there must b a way! i mean its an O'level ques. I hav my cambridge ppr tmmrw, & im badly stuckk. :/

myininaya (myininaya):

have you ever seen a problem like this one? can i find an example somewhere?

OpenStudy (saifoo.khan):

umm, my scanner isnt working now othrwise i should hav scannd & sndd u the example. Its in my book,

myininaya (myininaya):

is it too long to type in latex form?

myininaya (myininaya):

or you don't even have to use latex im sure i can make it out

OpenStudy (saifoo.khan):

:/

OpenStudy (saifoo.khan):

thnks for your Help.. =)

myininaya (myininaya):

maybe we can do something with pascals's triangle one sec

OpenStudy (saifoo.khan):

Yeah, its the same thing.

OpenStudy (anonymous):

the coefficient of the x term?

OpenStudy (saifoo.khan):

Yes. the ques says: Find the coefficient of x in the expansion

OpenStudy (anonymous):

if that is all you want it is -320

myininaya (myininaya):

satellite how did you do it so fast

OpenStudy (saifoo.khan):

Wow!

OpenStudy (saifoo.khan):

no, not wow.

myininaya (myininaya):

is that right khan?

OpenStudy (saifoo.khan):

its somrthing –960 + 320 = –640

OpenStudy (saifoo.khan):

this is the final answer

OpenStudy (saifoo.khan):

satelite??? HOWW?

myininaya (myininaya):

the way im doing it is too long lol

OpenStudy (anonymous):

do u want the highest index of the variable????

OpenStudy (anonymous):

ok here goes.

OpenStudy (anonymous):

\[(2-5x)^6 = 2^6 + stuff\]

OpenStudy (anonymous):

if yes then highest index will b 16

OpenStudy (anonymous):

and the x term in \[(1+x)^{10}=10x\]

OpenStudy (anonymous):

so one way to get an x is \[2^6\times 10x=640x\]

OpenStudy (anonymous):

the other way to get an x term is \[(1+x)^{10}=1+stuff\]

OpenStudy (anonymous):

and the x term in \[(2-5x)^6\] is \[6\times (-5x)^2\times 2^5\]

OpenStudy (anonymous):

binomial explansion yes? it is \[6ab^5\] where a = -5x and b = 2

myininaya (myininaya):

(-5x)^1*?

OpenStudy (anonymous):

i just put that in there to make it match up. the term is \[6ab^5\]

OpenStudy (anonymous):

so we get \[6\times (-5) \times 2^5=-960\]\]

OpenStudy (anonymous):

combining like terms, as we say in the math biz, gives \[-960x+640x=-320x\]

OpenStudy (anonymous):

k?

OpenStudy (anonymous):

oops yes myininaya i made a typo. it should be \[6(-5x)^12^5\] sorry

OpenStudy (anonymous):

it is the x term so certainly not \[(-5x)^2\] like i wrote. answer is right tho

myininaya (myininaya):

:)

OpenStudy (anonymous):

good thing they only asked for the x term. imagine what you would have to do if you wanted the x^2 term!

myininaya (myininaya):

i like to here you made a mistake it makes me feel good lol

OpenStudy (anonymous):

answer was right, so that only counts as a typo!

myininaya (myininaya):

no im putting this down in my scorebook verus you i think i might be winning now

myininaya (myininaya):

gj satellite on the problem

OpenStudy (anonymous):

no! answer right = typo. answer wrong = mistake! thnx btw you have way more fans than i do

myininaya (myininaya):

you have more medals i guess they just pretend to like me

OpenStudy (anonymous):

now i think it is beer o'clock.

myininaya (myininaya):

later

OpenStudy (saifoo.khan):

thnks Guys! =)